OpenStudy (anonymous):

How do you simplify (7-6i)(-8+3i)?

OpenStudy (jdoe0001):

multiply them :)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} (7-6i)(-8+3i) \implies& 7 \times (-8+3i) \implies -56+21i\\ \implies& -6i \times (-8+3i) \implies 48i-18i^2\\\quad \\\quad \\ \end{array}\\ \bf (7-6i)(-8+3i)\implies -56+21i+48i-18i^2\) add like-terms and expand any other ones